OpenStudy (word2):

Find the slope of the line containing the two points (1,-1) and (-5,-3). -1 -1/3 1/3 2 3

OpenStudy (word2):

my answer is -1/3 , am i right

OpenStudy (word2):

@agent0smith

OpenStudy (agent0smith):

If you show your work, I'll check it

OpenStudy (zzr0ck3r):

Part of learning math is knowing when you have done something correct and trusting that. So if you made all of the correct steps i.e. \(m=\dfrac{y_2-y_1}{x_2-x_1}\) and you got \(m=-\frac{1}{3}\) then you should trust it.

OpenStudy (word2):

thank you , but i got it wrong , lol the answer is in fact 1/3